Solve
(x+2)^2 = 7
Thanks for any help.

Answer by Nate(3500)

(x + 2)^2 = 7
sqrt((x + 2)^2) = +- sqrt(7)
x + 2 = +- sqrt(7)
x = -2 +- sqrt(7)